OpenStudy (gina):

how about this?,,,,xyy'=1-x^2

OpenStudy (anonymous):

\[xy.\frac{dy}{dx}=1-x^2\]This is again separable. Divide both sides by x and 'multiply' both sides by dx to get\[y.dy=\frac{1-x^2}{x}dx=(\frac{1}{x}-x)dx\]Integrate both sides:\[\frac{y^2}{2}=\ln(x)-\frac{x^2}{2}+c_1 \rightarrow y^2=2\ln(x)-x^2+c_2\]where c_1 and c_2 are just constants.

OpenStudy (gina):

thank u very much, sir

OpenStudy (anonymous):

You're welcome.